a^3-x * b^5x=a^x+5 * b^3x,find the value of x log(b/a)
Ans log a
Please give the solution

Grade:7

1 Answers

Sher Mohammad IIT Delhi
9 years ago
a3-xb5x=ax+5b3x
log(a3-xb5x)=log(ax+5b3x)
log(a3-x)+log(b5x)=log(ax+5)+log(b3x)
now, arrange term, gives
xlog(b/a)=loga